OpenStudy (anonymous):

Need help determining whether this series converges or not: \[\sum_{1}^{\infty}\frac{ 2n! }{ n!^2 }\] My thoughts are this is a candidate for ratio test, so I rewrote it \[\left| \frac{ 2(n+1)! }{ ((n+1)!)^2 } \right| * \left| \frac{ n!^2 }{ 2n! } \right|\] But I do not know how to simplify this expression.

OpenStudy (turingtest):

\[\frac{(n+1)!}{n!}=n+1\]did you know that?

OpenStudy (anonymous):

2(n+1)/2(n+1)

OpenStudy (turingtest):

don't drop the square on the other one...

OpenStudy (anonymous):

Wait thats wrong again, sorry, doing this in my head.

OpenStudy (anonymous):

2(n+1)/2(n+1)^2

OpenStudy (turingtest):

yes, and the 2's cancel...

OpenStudy (anonymous):

1/(n+1) so by the ratio test, if L<1 it converges?

OpenStudy (turingtest):

correct :)

OpenStudy (anonymous):

Really appreciate you working with me! Thanks

OpenStudy (turingtest):

very welcome!